OpenStudy (cheesymcdouble):

Explain the distance formula. then use it to calculate the distance between A(1,1) and B(7,-7). Sorry I suck at this stuff.

pooja195 (pooja195):

Well do you know what the distance formula is ?

pooja195 (pooja195):

Hint : DISTANCE

OpenStudy (cheesymcdouble):

oh ok

pooja195 (pooja195):

ok lets do that part later this is the distance formula \[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] you are given 2 coordinates A(1,1) and B(7,-7). (x1,y1)(x2,y2)

pooja195 (pooja195):

When you plug in the given coordinates you get this... \[\huge d=\sqrt{(7-1)^2+(-7-1)^2}\] Solve from here. :)

pooja195 (pooja195):

If you're stuck i can help you out! Just let me know! :)

OpenStudy (anonymous):



OpenStudy (cheesymcdouble):

sorry i am i really do suck at math...

OpenStudy (anonymous):

so it's 10

OpenStudy (anonymous):

My way is easier

pooja195 (pooja195):

@yomamabf they want you to use the distance formula read the directions. Congrats on giving the answer.

OpenStudy (anonymous):

oh i alwys thought my way was the distance formula we didn't even learn the real formula

OpenStudy (cheesymcdouble):

thank you so much @yomamabf sorry her way IS easier @pooja195

pooja195 (pooja195):

@cheesymcdouble Read the directions that were provided they want the distance formula not her way.

pooja195 (pooja195):

@yomamabf we need to do what the directions say.

OpenStudy (cheesymcdouble):

ok ok but it also wants the answer

pooja195 (pooja195):

you get the answer after you calculate this : \[\huge d=\sqrt{(7-1)^2+(-7-1)^2}\] Solve from here. :)

OpenStudy (cheesymcdouble):

alright thanks.

pooja195 (pooja195):

Also theres another part to your question it asks explain the distance formula are you fimiliar with what the formula is used for?

OpenStudy (cheesymcdouble):

wait what

pooja195 (pooja195):

This is why reading the directions is important.

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @cheesymcdouble Explain the distance formula. then use it to calculate the distance between A(1,1) and B(7,-7). Sorry I suck at this stuff. \(\color{#0cbb34}{\text{End of Quote}}\) Explain the distance formula

pooja195 (pooja195):

The distance formula is used to figure out the distance between 2 points. :)

OpenStudy (cheesymcdouble):

ok thanks

OpenStudy (anonymous):

fwiw both @pooja195 and @yomamabf applied the distance formula. The distance formula is just the specific application of the Pythagorean theorem when you have two points and want to know the distance between them. One method uses a formula, the other draws it out. Use the one that works best for you.

OpenStudy (cheesymcdouble):

ya the drawing one actually helps me best.

pooja195 (pooja195):

@peachpi but they want the asker to use the distance formula .-.

pooja195 (pooja195):

Oh nvm lol i thought they wanted the physical formula and not the drawing...

OpenStudy (anonymous):

yeah. it's the same either way. \(x_2-x_1=6\) and \(y_2-y_1=8\)

OpenStudy (cheesymcdouble):

ok thanks guys

OpenStudy (anonymous):

lol

OpenStudy (cheesymcdouble):